### v1

# Effective bias expansion for circumventing 21 cm foregrounds

Abstract
The 21 cm line of neutral hydrogen is a promising probe of the Epoch of Reionization (EoR) but suffers from contamination by spectrally smooth foregrounds, which obscure the large-scale signal along the line of sight. We explore the possibility of indirectly extracting the 21 cm signal in the presence of foregrounds by taking advantage of nonlinear couplings between large and small scales, both at the level of the density field and a biased tracer of that field. When combined with effective field theory techniques, bias expansions allow for a systematic treatment of nonlinear mode-couplings in the map between the underlying density field and the 21 cm field. We apply an effective bias expansion to density information generated with 21cmFAST using a variety of assumptions about which information can be recovered by different surveys and density estimation techniques. We find that the 21 cm signal produced by our effective bias expansion, combined with our least optimistic foreground assumptions, produces $\sim 30\%$ cross-correlations with the true 21 cm field. Providing complementary density information from high-redshift galaxy surveys yields a cross-correlation of 50-70%. The techniques presented here may be combined with foreground mitigation strategies in order to improve the recovery of the large-scale 21 cm signal.
I Introduction
The time from the birth of the first stars to the end of the Epoch of Reionization (EoR) marks an important era for the astrophysics of galaxy formation and assembly, as well as for the underlying growth of cosmological structure Loeb and Barkana (2001); Zaroubi (2013); Robertson (2022); McQuinn et al. (2007); Mesinger (2016); Liu et al. (2016). The redshifted 21 cm line from neutral hydrogen is a direct probe of this period of cosmic time Furlanetto et al. (2006); Pritchard and Loeb (2012), which is being actively targeted by many existing and proposed radio observatories Monsalve et al. (2017); Nambissan T. et al. (2021); Philip et al. (2019); Parsons et al. (2010); van Haarlem et al. (2013); Zarka et al. (2012); Tingay et al. (2013); Wayth et al. (2018); Gupta et al. (2017); DeBoer et al. (2017); Berkhout et al. (2024); Eastwood et al. (2019); Koopmans et al. (2015). In particular, radio interferometers have already set stringent upper bounds on the 21 cm power spectrum (Parsons et al., 2014; Dillon et al., 2014, 2015; Ewall-Wice et al., 2016; Beardsley et al., 2016; Patil et al., 2017; Li et al., 2019; Barry et al., 2019; Eastwood et al., 2019; Gehlot et al., 2019; Kolopanis et al., 2019; Trott et al., 2020; Mertens et al., 2020; Garsden et al., 2021; Yoshiura et al., 2021; Abdurashidova et al., 2022a; HERA Collaboration et al., 2023; Wilensky et al., 2023; Munshi et al., 2024; Acharya et al., 2024; Mertens et al., 2025; Nunhokee et al., 2025), which have been used to constrain the properties of the intergalactic medium (IGM) at high redshifts Abdurashidova et al. (2022b); HERA Collaboration et al. (2023); Ghara et al. (2025); Nunhokee et al. (2025).
A major ongoing issue in 21 cm cosmology is that the signal is highly contaminated by foreground emission that is many orders of magnitude brighter (see Ref. Liu and Shaw (2020) for a review). These foregrounds, which include Galactic synchrotron radiation, are expected to be spectrally smooth and will therefore affect large-scale modes of the 21 cm signal along the line of sight. Additionally, the chromatic response of interferometers can cause the foregrounds to contaminate other Fourier modes. Decomposing each wavevector into $\mathbf{k}\equiv(\mathbf{k}_{\perp},k_{\parallel})$, where $k_{\parallel}$ denotes the Fourier modes parallel to the line of sight and $\mathbf{k}_{\perp}$ represents the Fourier modes in the plane perpendicular to the line of sight, then in $_{\perp}-kk_{\parallel}$ space, the region where foreground contamination occurs is referred to as the “foreground wedge” due to its shape Datta et al. (2010); Morales et al. (2012); Parsons et al. (2012); Thyagarajan et al. (2015); Liu et al. (2014a, b); Munshi et al. (2025).
It is difficult to completely subtract foregrounds Datta et al. (2010); Liu and Tegmark (2011); Chapman et al. (2012); Ghosh et al. (2015); Ewall-Wice et al. (2021); Mertens et al. (2018); Wang et al. (2024) in the wedge due to uncertainties in modeling both the foreground emission and the chromatic response of the instrument Barry et al. (2016); Patil et al. (2016); Ewall-Wice et al. (2017); Byrne et al. (2019); Mouri Sardarabadi and Koopmans (2019); Neben et al. (2016); Joseph et al. (2018); Ansah-Narh et al. (2018); Zheng et al. (2017); Orosz et al. (2019); Joseph et al. (2020); Choudhuri et al. (2021); Kim et al. (2022); Ewall-Wice et al. (2016); Kern et al. (2019); Ung et al. (2020); Josaitis et al. (2022); Rath et al. (2024); Chen et al. (2025). Therefore, these modes are often discarded, with many analyses of 21 cm data focusing on evading foreground-contaminated modes Vedantham et al. (2012); Trott et al. (2012); Hazelton et al. (2013); Thyagarajan et al. (2013); Liu et al. (2014). This effectively renders a large number of cosmological modes inaccessible to current experiments, significantly reducing their sensitivity. It also limits the possibility of correlating the 21 cm signal with complementary tracers of the matter field from high-redshift galaxy surveys or line intensity mapping experiments La Plante et al. (2023); Gagnon-Hartman et al. (2025); Fronenberg and Liu (2024).
As with any biased tracer of the matter field, there is non-linear mode-coupling between large-scale and small-scale modes of the 21 cm field. In principle, it is therefore possible to take advantage of this mode-coupling to reconstruct large-scale information from uncontaminated measurements of small-scale modes. This type of reconstruction has been studied in the context of matter density and halo fields Zhu et al. (2016, 2022); Zang et al. (2024), as well as in 21 cm cosmology Zhu et al. (2018); Karaçaylı and Padmanabhan (2019); Darwish et al. (2021), although usually at low redshifts of $z\lesssim 5$. Past studies have also applied machine learning methods to reconstruct the foreground-obscured 21 cm modes Gagnon-Hartman et al. (2021); Kennedy et al. (2024); Sabti et al. (2025); Thélie et al. (2025). On length scales where 21 cm fluctuations are only mildly nonlinear, one can use techniques inspired by effective field theory (EFT) to systematically treat these nonlinearities McQuinn and D’Aloisio (2018); Qin et al. (2022). For instance, it was recently shown that EFT methods applied to small-volume (high-$k$) hydrodynamical simulations of the 21 cm field can accurately predict larger volumes Qin et al. (2025).
Figure 1: The steps involved in our procedure. Starting at the left, radio telescopes will measure high-$k$ 21 cm modes (step 1). Since the 21 cm signal is a biased matter tracer, one can infer the underlying matter density field information, which will also mostly be at high $k_{\parallel}$ values (step 2). We can then take advantage of nonlinear couplings to reconstruct the missing density modes (step 3) and then use the bias expansion to translate the density fluctuations back into 21 cm fluctuations (step 4).
In this work, we further explore the predictive power that comes from the coupling between modes with different Fourier shapes in order to see how well we can recover foreground-obscured 21 cm modes. To test the robustness of the techniques used here, we make several different assumptions about which 21 cm modes are contaminated by foregrounds and what underlying density information is available. We also assess the impact of supplemental density information inferred from other probes in different parts of Fourier space, such as galaxy surveys, e.g. by the Nancy Grace Roman Space Telescope Spergel et al. (2015).
The rest of this paper is organized as follows. In Section II, we provide a conceptual overview of how we reconstruct missing modes and describe the various assumptions we make about what density information is available. In Section III, we briefly review the effective 21 cm bias expansion of Refs. McQuinn and D’Aloisio (2018); Qin et al. (2022) and describe our methods for reconstructing missing density modes. In Section V, we present the results of applying the bias expansion to different combinations of density information, and also demonstrate the complementarity of measurements from galaxy surveys. We conclude in Section VI. Throughout this article, we use cosmological parameters from Planck 2018 Aghanim et al. (2020).
II Conceptual Overview
The pipeline for recovering the foreground-contaminated 21 cm signal, as schematically illustrated in Fig. 1, consists of four main steps:
  1. 1.
    Experiments such as HERA will measure the 21 cm signal outside of the foreground wedge.
  2. 2.
    From these 21 cm measurements, one can extract density information outside of the foreground wedge. With this density information and the 21 cm modes from step 1, one can infer the bias parameters describing their connection.
  3. 3.
    One can then reconstruct the density modes inside the foreground region by taking advantage of nonlinear mode-coupling in the matter density field.
  4. 4.
    Finally, with the bias parameters and linear density field, it is possible to map the density field back to 21 cm field using an effective bias expansion, thus obtaining 21 cm modes within foregrounds. We compare the predicted 21 cm modes to signals generated with 21cmFAST to validate this approach.
While the first step poses a challenging data analysis problem, the second step in this pipeline is highly nontrivial from a theoretical perspective. In principle, since the 21 cm radiation is a biased tracer of the density field, one should be able to express the 21 cm field in terms of an expansion in powers of the density field, described in Section III.1. This expansion could then be inverted in order to obtain the density fluctuations from the 21 cm fluctuations. However, in practice, nonlinear contributions make the information relatively difficult to extract Feng et al. (2018). These challenges have been extensively studied in the context of galaxy surveys, particularly for the purpose of reconstructing the linear density field and baryon acoustic oscillations Frisch et al. (2002); Brenier et al. (2003); Eisenstein et al. (2007); Jasche and Wandelt (2013); Kitaura (2013); Tassev and Zaldarriaga (2012); Wang et al. (2013); Schmittfull et al. (2017); Seljak et al. (2017); Hada and Eisenstein (2018); Modi et al. (2018); Liu et al. (2021); Zhu et al. (2022); Zang et al. (2024); Chen and Padmanabhan (2024), and some limited progress has been made based on using 21 cm information as a starting point Modi et al. (2019); Darwish et al. (2021); Zhou and Mao (2024).
Devising an optimal strategy for extracting density information from 21 cm fluctuations would be worth its own dedicated study, and therefore we choose to focus primarily on steps 3 and 4 in this work. However, we note that Ref. Chen et al. (in preparation) has shown that step 2 is possible by adopting a gradient-based sampler to simultaneously infer the bias parameters and the underlying density field from the observed 21 cm modes outside the foreground wedge. Specifically, deep in the perturbative regime, the 21 cm field and density field are primarily related through the linear bias term which is local Qin et al. (2022, 2025). This indicates that the density modes that can be recovered correspond to the same modes where one has uncontaminated 21 cm information. This is corroborated by the findings of Ref. Chen et al. (in preparation), which showed explicitly that density modes outside the foreground wedge can be reliably inferred with up to $90\%$ accuracy on large scales (depending on the region of Fourier space) assuming a $\sim 3\sigma$ detection of those same individual 21 cm modes. As it turns out, the same approach of Ref. Chen et al. (in preparation) can also be used to reconstruct some of the large-scale density modes in the foreground wedge because their forward model includes nonlinear mode-coupling. These results and future improvements to step 2 can interface with the framework presented here in a straightforward way.
Motivated by uncertainties about which modes can be reliably inferred, as this is an active line of inquiry, we consider multiple representative scenarios in the following discussion. Starting from the least optimistic case and moving to more ambitious possibilities, we assume that the density modes initially recovered are:
  • the Fourier modes $\mathbf{k}$ outside of the foreground wedge, which are those with Liu et al. (2014)
    $|k_{\parallel}|\leq\frac{D_{\mathrm{c}}(z)H(z)}{c(1+z)}|\mathbf{k}_{\perp}|,$ (1)
    where $D_{\mathrm{c}}(z)$ is the co-moving line-of-sight distance to redshift $z$ Hogg (1999), $H(z)$ is the Hubble parameter as a function of redshift, and $c$ is the speed of light. We will refer to this as the “wedge” case and it corresponds to the most likely future scenario.
  • the modes $k_{\parallel}>k_{\mathrm{cut}}$, corresponding to a scenario where densely spaced or specifically arranged interferometer array configurations make it possible to use foreground mitigation strategies that may be able to recover wedge modes at high $k_{\perp}$ Murray and Trott (2018); MacKay et al. (in preparation). We also include this case since in our discussion because it enables a more direct comparison with results from the literature Darwish et al. (2021). We will refer to this as the “flat” case.
  • the modes $k>k_{\mathrm{cut}}$, which has been used in the literature to approximate the limitations of galaxy surveys Darwish et al. (2021). We also include this case as further motivation for studies of 21 cm foreground mitigation, to show how much more of the missing modes we can recover given additional small scale information. We will refer to this as the “isotropic” case.
We treat all three of these cases the same way in steps 3 and 4, with the only difference being the availability of modes entering into the reconstruction.
III Nonlinear mode-coupling and reconstruction
III.1 An effective bias expansion of the 21 cm signal
Here, we briefly review how one can predict the large-scale 21 cm signal from the underlying density field by applying the effective bias expansion described in Refs. McQuinn and D’Aloisio (2018); Qin et al. (2022), which corresponds to the last step in Fig. 1.
In linear perturbation theory, density perturbations independently evolve with scale factor $a$ as $\delta^{(1)}\sim D(a)$, regardless of the length scale. Here $D(a)$ is the linear growth factor, corresponding to $D(a)\sim a$ in a matter-dominated universe. However, as structure formation progresses, nonlinear contributions increasingly couple different Fourier modes. We define $k_{\mathrm{NL}}$ as the wavenumber at which nonlinear contributions become significant, i.e. $k<k_{\mathrm{NL}}$ or length scales greater than $\sim 1/k_{\mathrm{NL}}$ can be treated perturbatively. In standard perturbation theory, the nonlinear evolution of $\delta$ is parameterized as an order-by-order expansion. Using bold subscripts to denote Fourier-transformed quantities, e.g. $(\delta)_{\mathbf{k}}\equiv\int d^{3}\mathbf{x}\,\delta(\mathbf{x})e^{-i\mathbf{k}\cdot\mathbf{x}}$, this expansion is given by
$\delta_{\boldsymbol{k}}=\sum_{n=1}^{\infty}a^{n}\delta^{(n)}_{\boldsymbol{k}}$, (2)
where the higher-order terms are in turn parameterized as convolutions of the linear density field,
$\delta^{(n)}_{\boldsymbol{k}}$ $=\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}_{1}\dots\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}_{n}\,(2\pi)^{3}\delta^{D}\left(\boldsymbol{k}-\sum_{i=1}^{n}\boldsymbol{q}_{i}\right)$
$\times F_{n}(\boldsymbol{q}_{1},\dots,\boldsymbol{q}_{n})\delta^{(1)}_{\boldsymbol{q}_{1}}\dots\delta^{(1)}_{\boldsymbol{q}_{n}}$, (3)
where $d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}$ is shorthand for $d^{3}\mathbf{q}/(2\pi)^{3}$. The Dirac delta function $\delta^{D}$ enforces a relationship between the coupled modes that enter into the convolution kernels $F_{n}$, which have well-known recursion relations Goroff et al. (1986); Jain and Bertschinger (1994); Bernardeau et al. (2002).
The matter velocity field $\mathbf{v}$ can be decomposed into its divergence, $\theta=\nabla\cdot\mathbf{v}$, and curl, $\omega=\nabla\times\mathbf{v}$. The curl component decays like $\omega\sim 1/a$ in standard perturbation theory, so we neglect it. Meanwhile, in linear perturbation theory $\theta^{(1)}\sim\delta^{(1)}$. The divergence of the velocity field can then be expressed as an expansion similar to the density field,
$\theta_{\boldsymbol{k}}=aH\sum_{n=1}^{\infty}a^{n}\theta^{(n)}_{\boldsymbol{k}}$, (4)
where higher-order terms are again expressed as convolutions of the linear density field,
$\theta^{(n)}_{\boldsymbol{k}}$ $=\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}q_{1}\dots\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}q_{n}\,(2\pi)^{3}\delta^{D}\left(\boldsymbol{k}-\sum_{i=1}^{n}\boldsymbol{q}_{i}\right)$
$\times G_{n}(\boldsymbol{q}_{1},\dots,\boldsymbol{q}_{n})\delta^{(1)}_{\boldsymbol{q}_{1}}\dots\delta^{(1)}_{\boldsymbol{q}_{n}}$, (5)
and where the recursion relations for the $G_{n}$ convolution kernels are also given in Refs. Goroff et al. (1986); Jain and Bertschinger (1994); Bernardeau et al. (2002).
We then assume that fluctuations in the 21 cm intensity field, $\delta_{21}$, trace the nonlinear matter density, and can thus be written as a local bias expansion in terms of $\delta$. We include all operators respecting rotational and Galilean invariance up to second order in $\delta$,
$(\delta_{21})_{\boldsymbol{k}}=$ $b_{1}\delta_{\boldsymbol{k}}-b_{\nabla^{2}}\left(\frac{k}{k_{\mathrm{NL}}}\right)^{2}\delta_{\boldsymbol{k}}$
$\quad+b_{2}\left(\delta^{2}\right)_{\boldsymbol{k}}+b_{\mathcal{G}2}(\mathcal{G}_{2})_{\boldsymbol{k}}+\cdots$ (6)
The tidal (Galileon) operator is defined in configuration space as
$\mathcal{G}_{2}=(\nabla_{i}\nabla_{j}\phi)(\nabla^{i}\nabla^{j}\phi)-(\nabla^{2}\phi)^{2},$ (7)
where $\phi$ is the gravitational potential. In addition to the linear and quadratic bias, the expansion in Eq. (6) includes $b_{\nabla^{2}}$, which captures local nonlinearities due to e.g. ionizing photons that travel long distances and is degenerate with leading-order EFT of LSS sound speed term Baldauf et al. (2015). We neglect the stochastic contribution to this expansion, as this term is negligible at the wavenumbers of interest McQuinn and D’Aloisio (2018).
The higher-order terms appearing in Eq. (6) involve integrating over nonlinear scales that are not well described using perturbative methods. For example, the term $\delta^{2}$ which appears as a product in configuration space is written in Fourier space as a convolution,
$\left(\delta^{2}\right)_{\boldsymbol{k}}=\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\boldsymbol{q}\,\delta_{\boldsymbol{q}}\delta_{\boldsymbol{k-q}},$ (8)
where the integration includes UV contributions from large wavenumbers (i.e. small length scales) which are beyond the regime of perturbative control. However, the long-wavelength modes that we are concerned with should be largely insensitive to small-scale nonlinear modes, and we thus add counterterms to each operator that cancel the UV-sensitive contributions up to a certain order in a procedure analogous to renormalization (see Refs. Assassi et al. (2014) and Qin et al. (2022) for more details). For example, the first few terms of the renormalized $\delta^{2}$ operator, which we denote as $\delta^{2}]$, are given by
$\delta^{2}]=\delta^{2}-\sigma^{2}(\Lambda)-\frac{68}{21}\sigma^{2}(\Lambda)\delta,$ (9)
where $\Lambda$ is the small-scale cutoff for the perturbative theory, $\sigma^{2}(\Lambda)=\int_{0}^{\Lambda}\frac{dq}{2\pi^{2}}\,q^{2}P_{\mathrm{L}}(q)$ is the variance in the linear matter fluctuations for $q<\Lambda$, and $P_{\mathrm{L}}(q)$ is the linear matter power spectrum. The bias parameters must also shift to compensate for the newly added counterterms, and for a given parameter $b_{i}$, we write its renormalized counterpart as $b_{i}^{(R)}$.
Finally, we incorporate contributions from redshift space distortions (RSDs). Due to the peculiar velocities of 21 cm sources, $\boldsymbol{v}_{\mathrm{pec}}$, the inferred coordinate of a source $\boldsymbol{x}_{r}$ is related to its true coordinate $\boldsymbol{x}$ by
$\boldsymbol{x}_{r}=\boldsymbol{x}+\frac{\hat{\boldsymbol{n}}\cdot\boldsymbol{v}_{\mathrm{pec}}}{aH}\hat{\boldsymbol{n}}$, (10)
where $\hat{\boldsymbol{n}}$ is the line-of-sight direction, $a$ is the scale factor, and $H$ is the Hubble parameter. One can then express the density field in redshift space in terms of the real-space density using Eq. (10) and obtain
$(\delta_{r})_{\boldsymbol{k}}=$ $\delta_{\boldsymbol{k}}-i\frac{k_{\parallel}}{aH}(v_{\parallel})_{\boldsymbol{k}}-i\frac{k_{\parallel}}{aH}(\delta v_{\parallel})_{\boldsymbol{k}}$
$-\frac{1}{2}\left(\frac{k_{\parallel}}{aH}\right)^{2}\left(v_{\parallel}^{2}\right)_{\boldsymbol{k}}-\frac{1}{2}\left(\frac{k_{\parallel}}{aH}\right)^{2}\left(\delta v_{\parallel}^{2}\right)_{\boldsymbol{k}}$
$+\frac{i}{6}\left(\frac{k_{\parallel}}{aH}\right)^{3}\left(v_{\parallel}^{3}\right)_{\boldsymbol{k}}+\cdots,$ (11)
where we have defined $v_{\parallel}=\hat{\boldsymbol{n}}\cdot\boldsymbol{v}_{\mathrm{pec}}$ and $k_{\parallel}=\hat{\boldsymbol{n}}\cdot\boldsymbol{k}$, and Taylor expanded the expression by treating $k_{\parallel}v_{\parallel}/aH$ as a small number. This expansion is justified since, on the length scales that we are interested in, the peculiar velocities of sources are a small effect compared to the Hubble flow. The velocity fields can be calculated from the linear density field using Eqs. (4) and (5). This RSD expansion then contains new operators such as $v_{\parallel}^{2}$ which must also be renormalized.
Combining Eqs. (6) and (11) and renormalizing all UV-sensitive operators, the relationship between the 21 cm field and linear density field is thus given by
$(\delta_{21,r})_{\boldsymbol{k}}=$ $~b_{1}^{(R)}\delta_{\boldsymbol{k}}-b_{\nabla^{2}}k^{2}\delta_{\boldsymbol{k}}+b_{2}^{(R)}\left[\delta^{2}\right]_{\boldsymbol{k}}+b_{G2}^{(R)}(\mathcal{G}_{2})_{\boldsymbol{k}}$
$-i\frac{k_{\parallel}}{aH}\left[(v_{\parallel})_{\boldsymbol{k}}+b_{1}(\delta v_{\parallel})_{\boldsymbol{k}}-b_{\nabla^{2}}k^{2}(\delta v_{\parallel})_{\boldsymbol{k}}\right]$
$-\frac{1}{2}\left(\frac{k_{\parallel}}{aH}\right)^{2}\left[v_{\parallel}^{2}\right]_{\boldsymbol{k}}+\cdots$ (12)
This model has been shown to reproduce the 21 cm power spectrum from hydrodynamical simulations with errors at the percent level, and the field-level signal with errors at the $\mathcal{O}(10\%)$ level Qin et al. (2022).
Since, throughout the text, we assume that both density and 21 cm measurements are available outside of some “wedge”-shaped, “flat”, or “isotropic” region, we fit the renormalized bias coefficients using Eq. (12) only over the available wavenumbers in each scenario. We perform a field-level fit for the coefficients by minimizing the function
$\mathcal{A}=\sum_{\mathbf{k}}\frac{|(\delta_{\mathrm{sim}})_{\mathbf{k}}-(\delta_{\mathrm{EFT}})_{\mathbf{k}}|^{2}}{V},$ (13)
where $\delta_{\mathrm{sim}}$ is the 21 cm fluctuations generated from e.g. 21cmFAST, $\delta_{\mathrm{EFT}}$ is the 21 cm signal predicted from the underlying density field using e.g. Eq. (12), and $V$ is the volume of the simulation Qin et al. (2022).
III.2 Reconstruction of missing density modes
For step 3 in Fig. 1, we use the procedure from Ref. Darwish et al. (2021) as one example for reconstructing missing modes at the level of the density field. Here, we summarize the main steps for constructing the estimator; more details about how to derive the following expressions can be found in Ref. Darwish et al. (2021).
In linear perturbation theory, different modes of an initially Gaussian density field evolve independently of one another; however, nonlinear gravitational evolution couples modes at different length scales. For example, the nonlinear density field can be written in terms of the linear matter overdensity $\delta^{(1)}$ as
$(\delta_{\mathrm{NL}})_{\mathbf{k}}=\delta^{(1)}_{\mathbf{k}}+\sum_{\alpha}\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}\,F_{\alpha}(\mathbf{q},\mathbf{k-q})\delta^{(1)}_{\mathbf{q}}\delta^{(1)}_{\mathbf{q-k}}$$\cdots$, (14)
where $\alpha$ runs over distinct types of quadratic couplings and $F_{\alpha}$ represents the corresponding kernel. The kernels we examine in this study correspond to (1) the “growth” coupling, which can be thought of as the enhanced or suppressed growth of a small-scale mode due to being embedded in a larger-wavelength over/underdensity,
$F_{G}(\mathbf{k}_{1},\mathbf{k}_{2})=\frac{17}{21};$ (15)
(2) the “shift” coupling, which can be thought of as the displacement of a mode due to the presence of another nearby over/underdensity,
$F_{S}(\mathbf{k}_{1},\mathbf{k}_{2})=\frac{1}{2}\left(\frac{1}{k_{1}^{2}}+\frac{1}{k_{2}^{2}}\right)\mathbf{k}_{1}\cdot\mathbf{k}_{2};$ (16)
and (3) the “tidal” coupling, which can be thought of as the tidal force on an overdensity/underdensity from being embedded in a varying gravitational field,
$F_{T}(\mathbf{k}_{1},\mathbf{k}_{2})=\frac{2}{7}\left[\frac{(\mathbf{k}_{1}\cdot\mathbf{k}_{2})^{2}}{k_{1}^{2}k_{2}^{2}}-\frac{1}{3}\right].$ (17)
At second order in standard perturbation theory, the kernel $F_{2}$ appearing in Eq. (3) is the sum of all three kernels listed above Jain and Bertschinger (1994); Ma and Bertschinger (1995). These couplings also arise in configuration-space expansions describing biased tracers of matter. For example, a quadratic bias term $\delta^{2}$ will have a kernel corresponding to the growth coupling. Similarly, the Fourier-transformed Galileon operator defined in Eq. (7) will have a kernel corresponding to the tidal coupling. Given density information in some part of Fourier space, we can use these mode couplings to reconstruct missing wavenumbers.
Given the nonlinear density field, the minimum variance unbiased estimator for the linear density field is given by Darwish et al. (2021)
$\hat{\Delta}_{\alpha,\mathbf{K}}$ $=N_{\alpha}(\mathbf{K})$ (18)
$\times\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}\,\frac{f_{\alpha}(\mathbf{q},\mathbf{K}-\mathbf{q})}{2P_{\mathrm{NL}}(q)P_{\mathrm{NL}}(|\mathbf{K}-\mathbf{q}|)}(\delta_{\mathrm{NL}})_{\mathbf{q}}(\delta_{\mathrm{NL}})_{\mathbf{K}-\mathbf{q}}$
where the normalization factor is given by
$N_{\alpha}(\mathbf{K})=\left(\int d\hskip-0.80002pt\bar{}\hskip 1.00006pt^{3}\mathbf{q}\frac{f_{\alpha}(\mathbf{q},\mathbf{K}-\mathbf{q})^{2}}{2P_{\mathrm{NL}}(q)P_{\mathrm{NL}}(|\mathbf{K}-\mathbf{q}|)}\right)^{-1}$ (19)
and also represents noise on the estimator from cosmic variance and shot noise, $f_{\alpha}$ is related to the kernels described previously by
$f_{\alpha}(\mathbf{q}_{1},\mathbf{q}_{2})=2\left[F_{\alpha}(\mathbf{q}_{1}+\mathbf{q}_{2},-\mathbf{q}_{1})P_{\mathrm{L}}(q_{1})+(1\leftrightarrow 2)\right],$ (20)
$P_{\mathrm{NL}}(q)$ is the nonlinear matter power spectrum, and $P_{\mathrm{L}}(q)$ the linear power spectrum. The expression in Eq. (18) is derived assuming there is only a single mode-coupling, e.g. only one of the growth, shift, or tidal type couplings present. Other couplings not accounted for in the estimator become sources of noise or contamination; see Ref. Darwish et al. (2021) for details.
We have validated our calculation by checking that we can reproduce $N_{\alpha}(\mathbf{K})$ in Ref. Darwish et al. (2021) in the wavenumber range where our results overlap when we make the same assumptions. Moreover, in agreement with Ref. Darwish et al. (2021), we find that the growth-type coupling has the lowest noise of the three listed above. Therefore, for the rest of this work, we will use the growth coupling when including the step of density reconstruction.
In Ref. Darwish et al. (2021), the normalization for the estimator is derived by requiring that it be unbiased at the field level, e.g. the expectation value of the estimator is the true linear density field. As a result, the cross-spectrum of the estimator with $\delta^{(1)}$ is given by
$\langle\delta^{(1)}_{\mathbf{k}}\hat{\Delta}_{\alpha,\mathbf{k}^{\prime}}\rangle=(2\pi)^{3}\delta(\mathbf{k}-\mathbf{k}^{\prime})P_{\mathrm{L}}(k).$ (21)
This property is appropriate for applications such as cross-correlating the estimator with galaxy surveys. However, this also means that the auto-spectrum of the estimator is
$\langle\hat{\Delta}_{\alpha,\mathbf{k}}\hat{\Delta}_{\alpha,\mathbf{k}^{\prime}}\rangle=(2\pi)^{3}\delta(\mathbf{k}-\mathbf{k}^{\prime})N_{\alpha}(k).$ (22)
In this work, since we are more concerned with the power spectrum of the estimator, we require that the estimator be unbiased at the level of the power spectrum. We accomplish this by replacing the normalization $N_{\alpha}$ in Eq. (18) and instead fixing the amplitude of each Fourier mode to its value inferred from the average power spectrum so that $|\hat{\Delta}_{\alpha,\mathbf{K}}|^{2}=P_{\mathrm{NL}}(k)$, which can be obtained from a cosmological Boltzmann code like CAMB Lewis et al. (2000) or CLASS Lesgourgues (2011). While this procedure artificially suppresses cosmic variance, it ensures that we recover the correct matter power spectrum, and subsequently the correct 21 cm power spectrum after applying the bias expansion described in the previous Section.
Since the validity of the estimator in Eq. (18) relies on the linear and quadratic terms dominating Eq. (14), the field $\delta_{\mathrm{NL}}$ must be smoothed using some window function $W(k)$ in order to exclude the high-$k$ modes that are the most nonlinear. The estimator in Eq. (18) can then be written in terms of convolutions, and therefore calculated using the convolution theorem. For example, if we substitute Eq. (15) into Eq. (20) to obtain $f_{G}$ and use this to calculate the estimator, then the integral in Eq. (18) can be written as a convolution over the fields
$(\delta_{A})_{\mathbf{k}}=\frac{(\delta_{\mathrm{NL}}^{W})_{\mathbf{k}}}{P_{\mathrm{NL}}(k)},\quad(\delta_{B})_{\mathbf{k}}=\frac{(\delta_{\mathrm{NL}}^{W})_{\mathbf{k}}P_{\mathrm{L}}(k)}{P_{\mathrm{NL}}(k)},$ (23)
where $\delta_{\mathrm{NL}}^{W}$ denotes the smoothed matter field. Hence, we can apply the inverse Fourier transformation to each of these fields, multiply the configuration-space fields, and Fourier transform the result to obtain the estimator. For the shift and tidal kernels, which have somewhat more complicated expressions, one can expand Eq. (18) into a sum of many convolutions over quantities similar to those appearing in Eq. (23).
To summarize, we use the available density information (e.g. outside of foregrounds) to calculate the expressions in Eq. (23). From $(\delta_{A})_{\mathbf{k}}$ and $(\delta_{B})_{\mathbf{k}}$ we can then calculate $\hat{\Delta}_{\alpha,\mathbf{K}}$, which provides an estimate of the missing density information.
IV Setup and mock fields
In order to calculate the right-hand side of Eq. (12), one either needs knowledge of both the matter density field and velocity field, or the linear density field alone since it can be used to calculate the full density through Eq. (2) and the velocity through Eq. (4). In what follows, we assume that we only have information about the underlying density and not the velocity. Therefore, we must assume knowledge of the linear density field even if the output of step 2 includes nonlinear contributions (i.e. the full matter density field) in actuality. As we are primarily concerned with the mildly nonlinear regime, we expect any contamination from these spurious contributions to be small since it only appears at higher order.
In addition to the density information obtained under the foreground scenarios described in Section II, we demonstrate the impact of various assumptions about the linear density $\delta^{(1)}$:
  • $\delta^{(1)}=\delta_{\mathrm{m,obs}}\oplus 0^{\text{FG}}$: the matter density only outside of the foreground region, which we assume can be inferred from 21 cm measurements, per step 2 in Section II. In the absence of information about the density, we set $\delta^{(1)}=0$ for modes inside the foreground-contaminated region before substituting into the effective bias expansion. This shows how well we can recover the 21 cm signal in the presence of foregrounds using only the bias expansion, in the absence of any density reconstruction. Fourier modes of $\delta_{21}$ within the foreground region only receive nonlinear contributions in this case due to their coupling to modes outside the region.
  • $\delta^{(1)}=\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,recon}}^{\text{FG}}$: the combination of $\delta_{\mathrm{m,obs}}$ and density information inside the foreground region reconstructed using the method described in Section III.2.
  • $\delta^{(1)}=\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,gal}}$: the combination of $\delta_{\mathrm{m,obs}}$ from 21 cm observations and matter density fluctuations inferred from galaxy surveys, which we describe further in Sec. V.2 (including our assumptions about redshift uncertainties). For modes that are within the 21 cm foreground region and not covered by galaxy surveys, we again set $\delta^{(1)}=0$.
  • $\delta^{(1)}=\delta_{\mathrm{m,full}}$: the full matter density field, both inside and outside the foreground region. While this is an overly optimistic assumption, the purpose here is to show the most accurate 21 cm signal that could in principle be recovered using our perturbative methods.
Figure 2: Terms in the bias expansion when applied to the $\delta_{\mathrm{m,obs}}$ with density modes missing in the foreground wedge. The first row shows the cross-correlation of each term with the true 21 cm signal, and the second row shows the ratio of the power spectrum of each term to the 21 cm power spectrum from 21cmFAST. Within the foreground region, the only term with a significant amplitude is the quadratic bias ($\delta^{2}$), which shows a $\sim 30\%$ correlation with the 21 cm fluctuations and is the dominant contribution to the bias expansion.
Figure 3: Same as Fig. 2, but applying density reconstruction to obtain density modes within the foreground wedge before mapping to 21 cm with the bias expansion. Compared to Fig. 2, including density reconstruction improves the recovery of both the phase information and amplitude of the linear terms, $\delta$ and $k^{2}\delta$. Nonlinear terms such as $\delta^{2}$ and $\mathcal{G}_{2}$ have an enhanced amplitude compared to Fig. 2. The sum of the terms, shown in the rightmost panel, has a similar cross-correlation to the true 21 cm signal as in Fig. 2, since the nonlinear terms already exhibited a significant cross-correlation.
Foregrounds # modes # perturbative
Full 2,406,104 41,851
Isotropic 2,403,899 39,646
Flat 2,100,852 18,362
Wedge 105,765 3,122
Table 1: The number of modes available from the 21cmFAST box under each assumption about modes lost to foregrounds, as well as the number of perturbative modes ($k<k_{\mathrm{NL}}$) in each case. The nonlinear scale $k_{\mathrm{NL}}$ and the foreground wedge are calculated at $z=8.5$.
The mock 21 cm brightness temperature fields that we compare our methods against are generated using 21cmFAST Mesinger et al. (2011); Murray et al. (2020), a semi-numerical modeling tool capable of efficiently producing realizations of the universe during the Cosmic Dawn and the EoR. 21cmFAST utilizes second-order Lagrangian perturbation theory (2LPT) to evolve the initial conditions of the universe to the desired redshift, and uses the excursion-set formalism Furlanetto et al. (2004) to generate the process of reionization. Such a formalism assumes that reionization is driven by galaxies and uses a set of empirical scaling relations to describe the astrophysical properties of these galaxies. In this work, we assume a saturated spin temperature and ignore the effect of X-ray heating. We adopt the astrophysical parametrization and fiducial parameters described in Ref. Park et al. (2019) to model reionization, which was shown to describe high-redshift galaxy luminosity functions well (up to $z\sim 10$). For the purpose of testing our reconstruction methods, we generate a $(500\,\mathrm{cMpc})^{3}$ box on a $134^{3}$ grid at a redshift of $z=8.5$.
We use 21cmFAST because of its ability to quickly generate such large volumes, which is necessary for testing the reconstruction of large-scale modes. The methods studied in this work could also be tested against other analytic and semi-analytic methods that can generate large volumes, such as Zeus21 Muñoz (2023). In this work, we do not validate our method against hydrodynamical simulations, which in principle capture a richer description of the astrophysical complexities underlying reionization. However, if such large hydrodynamical simulations were available, we expect that our methods would also work well in that context, since the effective bias expansion is able to accurately describe nonlinear mode-coupling for the purpose of enlarging simulations Qin et al. (2022, 2025).
After generating the fields with 21cmFAST, we remove modes that we assume are contaminated by foregrounds using the cuts described in Section III.2. For the “flat” and “isotropic” foregrounds, we use $k_{\mathrm{cut}}=0.15$ Mpc-1. The number of modes left uncontaminated by foregrounds is given in the second column of Table 1; the maximum number of modes is $134^{3}=2,406,104$.
To evaluate the expression in Eq. (18), we first filter the nonlinear small scales from the tracer field. As above, we denote the smoothed tracer field by $\delta_{\mathrm{NL}}^{W}$. We calculate the wavenumber cutoff $k_{\mathrm{NL}}$ for this filter by smoothing $\delta_{\mathrm{NL}}$ with a Gaussian window function on increasingly large length scales until the largest overdensity is less than 0.8; the value of $k_{\mathrm{NL}}$ is not particularly sensitive to the choice of value that we use for largest overdensity or the choice of window function Qin et al. (2022). For the 21cmFAST realization used in this work, we find that the nonlinear scale is approximately $k_{\mathrm{NL}}=0.4$ cMpc-1 at z=8.5. The third column of Table 1 shows the number of uncontaminated modes that are in the perturbative regime, i.e. with $k<k_{\mathrm{NL}}$. To evaluate the total power spectra in Eq. (23), we use the CAMB Boltzmann code Lewis and Challinor (2011).
Figure 4: Same as Fig. 3, but now assuming the flat (top panel) and isotropic (bottom panel) foreground shapes. Both cases show improved accuracy in the cross-correlation and power spectrum in foreground-contaminated regions as compared to the case with wedge-shaped foregrounds shown in Fig. 3.
V Results
In this Section, we apply the methods described in Section II to fields generated using 21cmFAST. We use two metrics for determining the success of recovering the 21 cm signal. The first is the cross-correlation between a field and the true 21 cm signal, which is given by
$r_{21}(\mathbf{k})=\frac{P_{X,21}(\mathbf{k})}{\sqrt{P_{X}(\mathbf{k})P_{21}(\mathbf{k})}},$ (24)
where $X$ represents any field. Note from the normalization of this expression that $r_{21}$ is insensitive to the amplitude of the field $X$ and therefore only encodes information about how well the phase of each Fourier mode in $X$ matches the 21 cm field realization. To capture the amplitude information, we can instead look at the ratio of the power spectrum of the desired field, which does not contain phase information, to the power spectrum of the mock 21 cm signal, $P_{X}(\mathbf{k})/P_{21}(\mathbf{k})$.
V.1 Individual terms
We first compare individual terms in the bias expansion to the mock 21 cm signal to determine which terms hold the greatest importance for an accurate reconstruction. We compute both $r_{21}$ and $P_{X}/P_{21}$ for each of the eight individual terms appearing in Eq. (12), including the best-fit bias coefficients and any other prefactors that multiply the RSD terms. We can then compare these to the values of $r_{21}$ and $P_{X}/P_{21}$ obtained by taking the sum of all the terms in the bias expansion. Note that we take this sum before computing the power spectrum, since the power spectrum of a sum is not equal to the sum of individual power spectra.
Fig. 2 shows the result of applying the bias expansion assuming $\delta^{(1)}=\delta_{\mathrm{m,obs}}$ outside of the foreground wedge and $\delta^{(1)}=0$ inside the wedge, i.e. without performing any density reconstruction. The only terms that show any significant amount of cross-correlation with the true 21 cm field in the wedge are those corresponding to $\delta^{2}$ and $\mathcal{G}_{2}$, with typical $r_{21}$ values of $0.3$ and 0.2-0.2, respectively. Moreover, $\delta^{2}$ is the only term with a large enough amplitude inside the wedge to contribute significantly to the overall bias expansion, with all other terms comprising sub-percent level contributions to the power spectrum in the foreground wedge. It is perhaps surprising that there is any appreciable level of correlation with the 21 cm field inside the wedge given that we are assuming that the initial linear density is identically zero in that region of Fourier space. Terms in Eq. (12) that are linear in $\delta^{(1)}$, which lack mode coupling, are therefore greatly suppressed relative to nonlinear terms such as $\delta^{2}$ and $\mathcal{G}_{2}$, although we note that even $\delta=\delta^{(1)}+\delta^{(2)}+\ldots$ has nonlinear contributions appearing in the linear bias. Nevertheless, these nonlinear terms by themselves do have some predictive power in the foreground wedge.
Applying the estimator described in Section III.2 to reconstruct density information within foregrounds (i.e. substituting $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,recon}}^{\text{FG}}$ into Eq. (12)), we obtain the term-by-term cross correlations and power spectra shown in Fig. 3. The linear terms in the expansion now show improved cross-correlation in the wedge at the level of about $\pm 0.2$, and also exhibit an increase in the amplitude of their power spectra. Fully nonlinear terms like $\delta^{2}$ and $\mathcal{G}_{2}$ show a similar level of cross-correlation as in Fig. 2; however, their amplitude in the foreground wedge is now magnified, with $P_{\mathcal{G}_{2}}$ exceeding the true power spectrum by a factor of about 3.5 and $P_{\delta^{2}}$ exceeding the true power spectrum by a factor of about 35. Although we see an improvement in the signal recovery on a term-by-term basis (particularly in the linear terms) once we include reconstructed density modes, we find that the improvement is less significant with all the terms summed together. This is because the most-improved terms are intrinsically less correlated to the true 21 cm field, and thus their better reconstruction does not significantly improve overall accuracy.
In Fig. 4, we perform a similar analysis assuming different foreground shapes prior to performing density reconstruction. In these cases, we see that the cross-correlation across the first several terms is higher than for the wedge-shaped foreground region: for the “flat” foregrounds, the first five terms all show cross-correlations of $|r_{21}|\sim 0.5$$0.6$, and for the “isotropic” foregrounds, we see cross-correlations of $|r_{21}|\sim 0.6$$0.7$. Since the “flat” and “isotropic” foreground shapes include more modes in the perturbative regime, compared to the wedge-shaped case (see again Table 1), this leads to an improved reconstruction of modes within the foreground region, and therefore the cross-correlation for the linear terms is now comparable to that of the nonlinear terms. This is also evident in the fact that the “isotropic” scenario contains more perturbative modes than the “flat” case, and therefore shows improved cross-correlation over the flat foregrounds.
Summing all the individual terms to obtain the total 21 cm field, we find that the cross-correlation in the foreground region is higher than for the wedge-shaped foregrounds, and the recovered amplitude of the power spectrum is closer to that of the 21cmFAST power spectrum. While the recovered power spectrum in the presence of wedge-shaped foregrounds is too large by a factor of about 23, this improves to only a factor of 6.4 for flat-shaped foregrounds and 4.0 for isotropic foregrounds. This improvement can again be attributed to the fact that these foreground shapes mean that our procedure starts with more perturbative modes, and thus we have better information with which to accurately reconstruct the amplitude of modes in the foreground region, as well as calculate the contribution of nonlinear terms.
Figure 5: The 21 cm signal resulting from substituting the density field with different modifications into Eq. 12. The first row shows the cross-correlation with the true 21 cm signal, $r_{\mathrm{21}}$, and the second row shows the ratio of the power spectra of the two fields. For the last three columns, we assume $\delta_{\mathrm{m,obs}}$ are the density modes lying outside of a wedge-shaped foreground region. In order to facilitate comparison of the different input density fields, we use the bias coefficients fit using the full density field across all columns.
V.2 Adding density information from galaxy surveys
Galaxies are a biased tracer of the underlying matter density, and a large number of galaxies present during the EoR will soon be observed thanks to ground and space-based surveys Bouwens et al. (2015); Finkelstein et al. (2015); Bagley et al. (2017); Rieke et al. (2023); Arrabal Haro et al. (2023). However, the density field inferred from galaxies is only accurate for wavenumbers less than some value of $k_{\parallel}$, as uncertainties in the measured redshifts of the observed galaxies $\sigma_{z}$ translate to uncertainties in their line-of-sight distances. Here, we assume a galaxy survey can measure the density field below $k_{\parallel}\leq H(z)/c/\sigma_{z}$ La Plante et al. (2023). We note that as galaxy surveys have high angular resolution, this $k_{\parallel}$ threshold is constant across $k_{\perp}$. Typically, photometric redshift uncertainties are around $\sigma_{z}\sim 0.5$ Bouwens et al. (2015); Finkelstein (2016) for dropout galaxies and $\sigma_{z}\sim 0.1$ for galaxies selected with a narrow-band filter Zheng et al. (2017); Kikuta et al. (2023). With follow-up observations by spectroscopic instruments, these measurements can be significantly improved with uncertainties as low as e.g. $\sigma_{z}\sim 0.001$ for the grism spectroscopy that will be available for the High-Latitude Wide Area Survey from the Nancy Grace Roman Space Telescope Spergel et al. (2015); Observations Time Allocation Committee and Community Survey Definition Committees (2025). In this discussion, we will be conservative by assuming a relatively large redshift uncertainty of $\sigma_{z}\sim 0.05$, with the understanding that improvements to this will only increase the level of signal recovery. This redshift uncertainty corresponds to measuring $k_{\parallel}<0.11h$ Mpc-1 in the density field.
In Fig. 5, we show $r_{21}$ and $P_{\mathrm{total}}/P_{21}$ for the 21 cm signal calculated assuming wedge-shaped foregrounds using different assumptions about the underlying density field. In order to ensure that differences between each column are not merely due to changes in the best-fit bias parameters, in each panel we use the same bias parameters as determined from the full density field. The first column shows the result of substituting the full density field, $\delta_{\mathrm{m,full}}$, into Eq. (12). This is the best recovery of the 21 cm signal that we could ever obtain using the bias expansion. Here, we see that the cross-correlation is about 0.9 for wavenumbers within the perturbative regime. The amplitude of the power spectrum is also recovered with high accuracy and is within a factor of 1.5 of the true 21 cm power spectrum.
The middle two columns serve as a comparison of the total reconstructed 21 cm field in the wedge with and without performing density reconstruction in the wedge. Hence, they are quite similar to the final columns of Figs. 2 and Figs. 3, but as explained above are constructed using a slightly different set of bias parameters. As discussed in Section V.1, the cross-correlation is slightly better without first performing density reconstruction. In other words, even if we assume that the matter density in the foreground wedge is zero, which is clearly incorrect, the mode-coupling intrinsic to the EFT terms is enough to yield a cross-correlation at the level of $r\sim 0.3$. This behavior can be understood as arising from the relative contribution of the linear terms with and without in-wedge density reconstruction, as described in Section V.1: the slight degradation of the cross-correlation is due to the fact that the linear terms in the bias expansion are less correlated with the true 21 cm fluctuations than the nonlinear terms. Additionally, without first performing density reconstruction, the reconstructed in-wedge 21 cm field underestimates the power spectrum by a factor of 8, due to the fact that we are missing contributions to the 21 cm fluctuations from many of the underlying density modes, whereas with density reconstruction the reconstructed 21 cm field overestimates the power spectrum by a factor of 90 (this is larger than what was reported in the previous section due to the different bias coefficients).
The last column of Fig. 5 shows the result of substituting $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,Roman}}$ into Eq. (12). In this case, for modes that are within the foreground wedge and also not covered by the galaxy survey, the recovered cross-correlation is at the level of 0.5, and the power spectrum is within a factor of a few of the true amplitude. This is a clear improvement over using either $\delta_{\mathrm{m,obs}}$ or $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,recon}}$. This highlights the complementarity of high redshift galaxy surveys to 21cm experiments, and motivates their joint use in recovering the underlying density field as well as reconstruction of 21 cm modes obscured by foregrounds.
VI Conclusion
Separating the information in long-wavelength modes along the line of sight from foregrounds is a longstanding problem of 21 cm cosmology. Bias expansions provide a map between density fluctuations and tracers such as 21 cm fluctuations, and may thus be used to indirectly recover 21 cm information within foregrounds. In this work, we have tested how well one can reconstruct the 21 cm fluctuations using the bias expansion under different scenarios about the available density information.
Assuming that we have a method of obtaining density information outside of the foreground wedge Chen et al. (in preparation), we find that simply applying our bias expansion to these modes already yields $\sim 30\%$ cross-correlations with the true 21 cm signal within the foreground region, due to non-linear couplings between small- and large-scale modes. Applying the density reconstruction method of Ref. Darwish et al. (2021) to obtain additional density information within the foreground wedge improves the recovery of linear contributions to the 21 cm signal within the foreground region. For the wedge-shaped foreground scenario, density reconstruction does not improve the overall cross-correlation. Density reconstruction applied to the “flat” and “isotropic” scenarios yields results that are much improved compared to the density-reconstructed “wedge” case, giving cross-correlations as high as $70\%$ and power spectra that are within a factor of a few of the true value—this motivates further study of independent methods for extracting small-scale density information.
For comparison, we also study the impact of including density information from galaxy surveys. This information is very complementary, as galaxy surveys tend to be sensitive to low values of $k_{\parallel}$ less than some value corresponding to redshift uncertainties, which are the most contaminated modes in 21 cm observations. We find that with the inclusion of galaxy survey information, the cross-correlation of recovered 21 cm modes with the true 21 cm signal can be as high as $50\%$ even for our least optimistic foreground assumptions.
From an information theory point of view, the reconstructed large-scale 21 cm modes do not add additional information at the field level, as they are determined solely from the small-scale modes. However, our results nonetheless demonstrate that the EFT is capable of capturing mode couplings generated by astrophysics-driven modeling frameworks such as 21cmFAST. Moreover, these reconstructed large-scale modes enable straightforward cross-correlation with other large-scale probes such as galaxy surveys without resorting to higher-order statistics such as the bispectrum. Our method also provides a potential route to create large-scale 21 cm templates at the field level to uncover foreground-contaminated modes in the data through a matched-filter approach Wang et al. (2024).
We emphasize that the methods presented here can build upon density reconstruction methods such as those presented in Ref. Darwish et al. (2021), but are conceptually different. While that work used measurements of biased tracers to reconstruct the matter density on scales obscured by foregrounds, the present work is concerned with whether one can use information about the density in small scales to obtain the 21 cm modes obscured by foregrounds. In this work, we have only explored the interface of the two methods by using the density estimator of Ref. Darwish et al. (2021) to supplement the small-scale density information assumed here. However, it may be instructive in the future to explore other ways in which these methods complement each other.
In summary, we have demonstrated that bias expansions combined with EFT-based treatments of nonlinear mode-coupling provide a useful starting point for recovering 21 cm information lost to foregrounds. Their utility is enhanced when combined with reliable extractions of the underlying density field within the foreground region, e.g. from galaxy surveys. There are a number of potential directions for improvement on the results presented here. For example, the density estimator of Ref. Darwish et al. (2021) that we apply throughout is derived using an unrenormalized bias expansion; it would be instructive to rederive this estimator using renormalized operators and including all of the appropriate counterterms. Moreover, an assumption that we make throughout this work is that one can extract density information outside of the 21 cm foreground wedge from 21 cm measurements due to the strength of the linear bias term. However, this assumption merits a detailed study in its own right and new techniques to extract this density information, such as those presented in Ref. Chen et al. (in preparation), could be combined with the techniques studied here in a straightforward way. We leave these as directions for future study.
Acknowledgments
It is a pleasure to thank Simon Foreman and Julian Muñoz for helpful discussions and comments on this paper. W.Q. was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064 and a grant from the Simons Foundation (Grant Number SFI-MPS-SFJ-00006250, W.Q.). K.-F.C. acknowledges support from the Mitacs Globalink Research Award, Taiwan Think Global Education Trust Scholarship, and the Taiwan Ministry of Education’s Government Scholarship to Study Abroad. K.S. acknowledges support from a Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant and from the Canada Research Chairs program. A.L. acknowledges support from an NSERC Discovery Grant and a Discovery Launch Supplement, and the William Dawson Scholarship at McGill. This analysis made use of Numpy Harris et al. (2020), Scipy Virtanen et al. (2020), Jupyter Kluyver et al. (2016), tqdm da Costa-Luis (2019), and Matplotlib Hunter (2007).